OpenStudy (anonymous):

How do you find the constant c when you've already found the taylor series expansion of lnx about c=2.

OpenStudy (anonymous):

c=2? I don't understand the question

OpenStudy (anonymous):

Is it the case you've expanded about the point c and you have something like, \[\sum_{}{}=\ln c + \frac{x-c}{c}-\frac{(x-c)^2}{2c^2}+...\]and you want to know what to do now?

OpenStudy (anonymous):

i.e. with ln(c)?